Saginaw Company is a garden products wholesale firm. In December, Saginaw Company expects to sell 26,000 bags of vegetable fertilizer at an average price of $5.6 per bag. Actual results are 26,700 bags sold at an average price of $5.5 per bag.

Required:

1. Calculate the sales price variance for December.
$

2. Calculate the sales volume variance for December.
$

3. Calculate the overall sales variance for December.
$

4. What if December sales were actually 25,800 bags calculate the new sales price variance?
$

The sales volume variance?
$

The overall sales variance?
$
Answer #1

1) Sales price variance = (5.6-5.5)*26700 = 2670 U

2) Sales volume variance = (26000-26700)*5.6 = 3920 F

3) Overall sales variance = (26000*5.6)-(26700*5.5) = 1250 F

4) Sales price variance = (5.6-5.5)*25800 = 2580 U

5) Sales volume variance = (26000-25800)*5.6 = 1120 U

6) Overall sales variance = (26000*5.6)-(25800*5.5) = 3700 U